# T2AcmrTempora-TLF

### v2

@@ 8 Ethics
  • Wang et al. (2022) Yizhong Wang, Swaroop Mishra, Pegah Alipoormolabashi, Yeganeh Kordi, Amirreza Mirzaei, Atharva Naik, Arjun Ashok, Arut Selvan Dhanasekaran, Anjana Arunkumar, David Stap, Eshaan Pathak, Giannis Karamanolakis, Haizhi Lai, Ishan Purohit, Ishani Mondal, Jacob Anderson, Kirby Kuznia, Krima Doshi, Kuntal Kumar Pal, and et al. 2022. Super-naturalinstructions: Generalization via declarative instructions on 1600+ nlp tasks. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 5085–5109, Abu Dhabi, United Arab Emirates. Association for Computational Linguistics.
@@
  • Yeshpanov et al. (2024) Rustem Yeshpanov, Pavel Efimov, Leonid Boytsov, Ardak Shalkarbayuli, and Pavel Braslavski. 2024. Kazqad: Kazakh open-domain question answering dataset. In arXiv preprint arXiv:2404.04487.
@@
  • Yeshpanov et al. (2022) Rustem Yeshpanov, Yerbolat Khassanov, and Huseyin Atakan Varol. 2022. KazNERD: Kazakh named entity recognition dataset. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 417–426, Marseille, France. European Language Resources Association.
  • Yeshpanov and Varol (2024) Rustem Yeshpanov and Huseyin Atakan Varol. 2024. Kazsandra: Kazakh sentiment analysis dataset of reviews and attitudes.
@@
  • Zhang et al. (2024) Meishan Zhang, Gongyao Jiang, Shuang Liu, Jing Chen, and Min Zhang. 2024. LLM-assisted data augmentation for Chinese dialogue-level dependency parsing. Computational Linguistics, 50(3):867–891.
@@
  • (42) Tianyi Zhang, Varsha Kishore, Felix Wu, Kilian Q Weinberger, and Yoav Artzi. Bertscore: Evaluating text generation with bert. In International Conference on Learning Representations.

### v3

@@ 8 Ethics
  • Y. Wang, S. Mishra, P. Alipoormolabashi, Y. Kordi, A. Mirzaei, A. Naik, A. Ashok, A. S. Dhanasekaran, A. Arunkumar, D. Stap, E. Pathak, G. Karamanolakis, H. Lai, I. Purohit, I. Mondal, J. Anderson, K. Kuznia, K. Doshi, K. K. Pal, and et al. (2022) Super-naturalinstructions: generalization via declarative instructions on 1600+ nlp tasks. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, Abu Dhabi, United Arab Emirates, pp. 5085–5109. Cited by: §2.1.
@@
  • R. Yeshpanov, P. Efimov, L. Boytsov, A. Shalkarbayuli, and P. Braslavski (2024) KazQAD: kazakh open-domain question answering dataset. In arXiv preprint arXiv:2404.04487, External Links: Link Cited by: §2.3.
@@
  • R. Yeshpanov, Y. Khassanov, and H. A. Varol (2022) KazNERD: Kazakh named entity recognition dataset. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, Marseille, France, pp. 417–426. External Links: Link Cited by: §1.
  • R. Yeshpanov and H. A. Varol (2024) KazSAnDRA: Kazakh sentiment analysis dataset of reviews and attitudes. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), N. Calzolari, M. Kan, V. Hoste, A. Lenci, S. Sakti, and N. Xue (Eds.), Torino, Italia, pp. 9657–9667. External Links: Link Cited by: §1.
@@
  • M. Zhang, G. Jiang, S. Liu, J. Chen, and M. Zhang (2024) LLM-assisted data augmentation for Chinese dialogue-level dependency parsing. Computational Linguistics 50 (3), pp. 867–891. External Links: Link, Document Cited by: §1.
@@
  • [43] T. Zhang, V. Kishore, F. Wu, K. Q. Weinberger, and Y. Artzi BERTScore: evaluating text generation with bert. In International Conference on Learning Representations, Cited by: §5.2.